INFRARED SPECTRA OF THE NE_{2}-N_{2}O, AR_{2}-N_{2}O TRIMERS

\underline{M. REZAEI}, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; K.H. MICHAELIAN, Natural Resources Canada, CanmetENERGY, 1 Oil Patch Drive, Suite A202 Devon, Alberta, T9G 1A8, Canada; A.R.W. McKELLAR, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

Spectra of the van der Waals trimers Ar_{2}-N_{2}O and Ne_{2}-N_{2}O are studied in the region of the N_{2}O ν_{1} fundamental (∼2224 cm^{−1}) using a tunable quantum cascade laser to probe a pulsed supersonic expansion from a slit jet nozzle. Improved data are also obtained for the dimers Ar-N_{2}O and Ne-N_{2}O, with the latter representing a significant improvement on the best previous results. As well, a feature in the spectrum is tentatively assigned as the Q-branch of Ar_{3}-N_{2}O. The observed vibrational shifts for Ne_{n}-N_{2}O are almost exactly linear for n = 0-2. However, for Ar_{n}-N_{2}O the n = 2 band origin is slightly blue-shifted compared to the linear prediction, and the n = 3 origin (if correct) is more significantly blue-shifted (by 0.09 cm^{−1}).